# Game-Theoretic Defenses for Robust Conformal Prediction Against Adversarial Attacks in Medical Imaging

### v1

@@ Abstract.
Adversarial attacks pose significant threats to the reliability and safety of deep learning models, especially in critical domains such as medical imaging. This paper introduces a novel framework that integrates conformal prediction with game-theoretic defensive strategies to enhance model robustness against both known and unknown adversarial perturbations. We address three primary research questions: constructing valid and efficient conformal prediction sets under known attacks (RQ1), ensuring coverage under unknown attacks through conservative thresholding (RQ2), and determining optimal defensive strategies within a zero-sum game framework (RQ3). Our methodology involves training specialized defensive models against specific attack types and employing maximum and minimum classifiers to aggregate defenses effectively. Extensive experiments conducted on the MedMNIST datasets—PathMNIST, OrganAMNIST, and TissueMNIST—demonstrate that our approach maintains high coverage guarantees while minimizing prediction set sizes. The game-theoretic analysis reveals that the optimal defensive strategy often converges to a singular robust model, outperforming uniform and simple strategies across all evaluated datasets. This work advances the state-of-the-art in uncertainty quantification and adversarial robustness, providing a reliable mechanism for deploying deep learning models in adversarial environments.
Rui Luo***ruiluo@cityu.edu.hk1, Jie Bao1486103897@qq.com2, Zhixin Zhouzhixin0825@gmail.com1, Chuangyin Dang§§§mecdang@cityu.edu.hk1
@@ 1. Introduction
Adversarial attacks [1] significantly undermine the reliability and safety of deep learning models, particularly in high-stakes domains such as medical diagnostics [2] and autonomous driving [3]. These attacks introduce subtle perturbations to input data, causing models to produce incorrect and potentially harmful predictions. In critical applications like healthcare and autonomous systems, ensuring both adversarial robustness and reliable uncertainty quantification is paramount. Conformal Prediction (CP) [4, 5, 6] offers a robust framework for uncertainty quantification by generating prediction sets that encompass the true label with a predefined confidence level [7].
@@
The interplay between adversarial robustness and conformal prediction [8, 9] is both crucial and underexplored. While extensive research has focused on enhancing adversarial robustness through various adversarial training methodologies [10, 11, 12, 13], the impact of these methods on the efficiency and reliability of conformal prediction sets remains inadequately understood. Specifically, prior studies have concentrated on adapting the scoring functions to attacks on either the test data or the calibration data, or both. However, these approaches often encounter validity issues, thereby compromising the reliability of CP sets under adversarial conditions.
This paper aims to bridge this knowledge gap by constructing robust conformal prediction sets for medical imaging datasets under diverse adversarial attack scenarios. We employ various attack types and partition the training data to develop specialized models resistant to each specific attack. Our objective is to formulate CP sets that uphold coverage guarantees while minimizing and optimizing prediction set sizes, even when subjected to both known and novel adversarial attacks.
To address the validity issues inherent in previous methodologies, we construct conservative prediction sets based on the largest estimated quantile derived from calibration data under different attacks. This approach ensures that the prediction sets maintain their validity guarantees despite adversarial perturbations. Central to our investigation is a pivotal question: given that validity is preserved, what constitutes the severest attack an adversary can orchestrate when they can choose any attack without the defender’s knowledge? Furthermore, how can the defender effectively mitigate such attacks under this adversarial scenario? Addressing these questions leads naturally to a zero-sum game formulation, where we consider both maximum and minimum classifiers to strategize optimal defenses against the most potent attacks.
By integrating game-theoretic principles with conformal prediction, this study endeavors to enhance the robustness and reliability of deep learning models in adversarial environments. Our approach not only maintains high coverage rates but also ensures minimal prediction set ambiguity, advancing the safety and efficacy of AI deployments in critical applications such as healthcare.
@@ 2.1. Adversarial Attack
However, poisoning attacks are more difficult to detect in deep learning models. Consequently, a wider range of methods has been developed to enhance model robustness against such attacks. For instance, improvements have been made by modifying the model architecture [29], designing robust loss functions [30, 31], and optimizing loss functions [32].
@@ 2.2. Conformal Prediction under Adversarial Attack
Uncertainty estimation is crucial for the robustness of deep learning models. Conformal Prediction (CP), introduced by Vovk et al. [7], provides distribution-free coverage guarantees but faces challenges when confronted with data poisoning and adversarial attacks. Gendler et al. [33] proposed RSCP, enhancing CP’s robustness through randomized smoothing by replacing the scoring function to defend against $\ell_{2}$-norm adversarial perturbations, though its formal guarantees remain limited. Yan et al. [34] introduced RSCP+, which adds a hyperparameter $\beta$ and Hoeffding bounds to correct Monte Carlo errors, while incorporating RCT and PTT techniques to improve efficiency. However, these modifications often result in overly conservative prediction sets and increased reliance on the holdout set.
@@
In contrast, Ghosh et al. [35] proposed PRCP, which approaches adversarial perturbations from a probabilistic standpoint by relaxing robustness guarantees under a predefined distribution. Cauchois et al. [36] addressed distributional shifts using an f-divergence constraint, though the resulting optimization is overly conservative [37]. Angelopoulos et al. [8] extended CP to control expected loss but did not provide algorithms for computing worst-case adversarial loss. Einbinder et al. [9] demonstrated that standard CP is robust to random label noise, though adversarial label perturbations were not explored. Furthermore, traditional defenses based on anomaly detection, loss function adjustments, and data clustering [15, 16, 17, 19, 22] struggle to safeguard CP’s unique properties, especially against sophisticated poisoning attacks targeting uncertainty mechanisms.
@@
Building on these limitations, Liu et al. [12] proposed an uncertainty reduction-based adversarial training method (AT-UR), which combines Beta-weighted loss and entropy minimization to enable models to maintain robustness while generating larger and more meaningful prediction sets. Jeary et al. [13] introduced Verifiably Robust Conformal Prediction (VRCP), which supports perturbations bounded by various norms (including $\ell_{1}$, $\ell_{2}$, and $\ell_{\infty}$) and regression tasks, using recent neural network verification techniques to restore coverage guarantees under adversarial attacks. Li et al. [11] developed a new class of black-box data poisoning attacks targeting Conformal Prediction (CP), where the attacker aims to manipulate prediction uncertainty rather than induce misclassification. They also proposed an optimization framework to defend against such attacks, demonstrating its effectiveness across a range of scenarios. Zargarbashi et al. [10] proposed using the cumulative distribution function (CDF) of smoothed scores to derive tighter upper bounds on worst-case score variations, leading to more efficient prediction sets while preserving robustness guarantees. Their method addresses both continuous and discrete/sparse data, offering guarantees for evasion (perturbing test inputs) and contamination attacks (perturbing calibration data). Scholten et al. [38] introduced Reliable Prediction Sets (RPS), which aggregate predictions from classifiers trained on different partitions of the training data and calibrated on disjoint subsets, making CP more resilient to data poisoning attacks.
@@
Recent studies have also highlighted CP’s resilience to label noise. Penso and Goldberger [39] introduced Noise-Robust Conformal Prediction (NR-CP), which estimates noise-free scores from noisy data, constructing smaller and more efficient prediction sets while maintaining coverage guarantees. NR-CP significantly outperforms other methods in noisy environments. Another study [9] explored the inherent robustness of CP, showing that even with noisy labels, CP tends to conservatively maintain coverage, although some adversarial noise conditions may compromise its reliability. Both studies emphasize CP’s resilience to label noise, a perturbation type that, while not intentionally malicious, introduces variability similar to adversarial attacks.
@@
In summary, our research complements and extends existing defense strategies by integrating game theory to model adversarial interactions and proposing novel mechanisms to fortify CP systems. This approach not only builds on the strengths of previous studies [35, 13, 11, 10] but also addresses their limitations, providing enhanced robustness and reliability for CP in high-stakes, adversarial environments.
@@ 3. Preliminary and Problem Setup
In this section, we establish the foundational notation and problem setup for evaluating and enhancing the robustness of Conformal Prediction (CP) systems against adversarial attacks. By defining the adversarial attack strategies, defensive models, conformal prediction framework, score functions, and metrics, and by formulating key research questions, we set the stage for developing and analyzing robust prediction mechanisms that maintain valid uncertainty estimates in the presence of deliberate perturbations. The notation used throughout the paper is given in Table 1.
@@ 3.3. Problem Formulation
To formalize these questions, consider the following setup: Let $f_{0}$ be the normal model trained on clean data $\{(x_{i},y_{i})\}_{i\in\mathcal{I}_{\textrm{train}}}$. An adversary with access to the true labels of the test dataset applies an attack function $g_{j}$ to generate perturbed inputs $x^{\prime}_{i}=g_{j}(x_{i},y_{i},f_{0})$ for $\{(x_{i},y_{i})\}_{i\in\mathcal{I}_{\textrm{test}}}$. Defensive models $\{f_{j}\}$ are trained on $\{(x^{\prime}_{i},y_{i})\}_{i\in\mathcal{I}_{\textrm{train}}}$ to resist specific attacks, where $x^{\prime}_{i}=g_{j}(x_{i},y_{i},f_{0})$, i.e., each $f_{j}$ corresponds to defending against attack $g_{j}$.
@@ 4.2. Addressing RQ2: Unknown Adversarial Attack
1:Training dataset $\{(x_{i},y_{i})\}_{i\in\mathcal{I}_{\textrm{train}}}$, Calibration dataset $\{(x_{i},y_{i})\}_{i\in\mathcal{I}_{\text{cal}}}$, Test dataset $\{x_{i}\}_{i\in\mathcal{I}_{\textrm{test}}}$, Set of potential attack functions $g_{j},j=1,\dots,m$, Coverage probability $1-\alpha$ 2:Prediction sets $\left(\Gamma(x_{i})\right),i\in\mathcal{I}_{\textrm{test}}$ 3:Based on the training method for $f_{k}$ outlined in Algorithm 2, we obtain the pre-trained defensive models $f_{k}$, $k=1,\dots,m$. $\triangleright$ Determine the largest quantile score across various attacks $g_{j}$ on the calibration set: 4:for each attack function $g_{j},j\in 1,\dots,m$ do       Apply attack $g_{j}$ to the calibration set to obtain:       $x^{\prime}_{i}=g_{j}(x_{i},y_{i},f_{0}),i\in\mathcal{I}_{\text{cal}}$.
      Use $f_{k}(x^{\prime}_{i})$ to compute non-conformity scores $\{s^{j}(x^{\prime}_{i},y_{i})\}_{i\in\mathcal{I}_{\text{cal}}}$.       Compute the quantile threshold as:       $q^{j}_{1-\alpha}$ as the $\lceil(1+|\mathcal{I}_{\text{cal}}|)(1-\alpha)\rceil$-th smallest score in ${s^{j}(x^{\prime}_{i},y_{i})}_{i\in\mathcal{I}_{\text{cal}}}$.
5:Determine the maximum quantile threshold $q_{1-\alpha}=\max_{j=1,\dots,m}q^{j}_{1-\alpha}$. $\triangleright$ Construct conservative prediction sets using the largest quantile score to ensure coverage: 6:for each test instance $x_{i}$, $i\in\mathcal{I}_{\textrm{test}}$ do       Construct the conservative prediction set: $\Gamma(x_{i})=\{y\in\mathcal{Y}\mid s(x_{i},y)\leq q_{1-\alpha}\}$
Algorithm 3 Constructing Conformal Prediction Sets under Unknown Attack (RQ2)

### v2

@@ Abstract.
Adversarial attacks pose significant threats to the robustness of deep learning models, especially in critical domains such as medical imaging. This paper introduces a framework that integrates conformal prediction with game-theoretic defensive strategies to counter both known and partially unknown adversarial perturbations. We address three research questions: constructing valid and efficient conformal prediction sets under known attacks (RQ1), ensuring coverage under unknown attacks through conservative thresholding (RQ2), and determining optimal defensive strategies within a zero-sum game framework (RQ3). In our methodology, specialized defensive models are trained against specific attack types and aggregated via maximum and minimum classifiers. Extensive experiments on the MedMNIST datasets—PathMNIST, OrganAMNIST, and TissueMNIST—show that our approach preserves high empirical coverage while minimizing prediction set sizes. The game-theoretic analysis indicates that the optimal defense strategy often converges to a singular robust model, outperforming simple uniform approaches. This work advances uncertainty quantification and adversarial robustness, providing a reliable mechanism for deploying deep learning models in adversarial environments.
Rui Luo***ruiluo@cityu.edu.hk1, Jie Bao1486103897@qq.com2, Zhixin Zhouzhixin0825@gmail.com3, Chuangyin Dang§§§mecdang@cityu.edu.hk1
@@ 1. Introduction
Adversarial attacks [1] significantly undermine the robustness of deep learning models, particularly in high-stakes domains such as medical diagnostics [2] and autonomous driving [3]. These attacks introduce subtle perturbations to input data that can lead to incorrect and potentially harmful predictions. In critical applications like healthcare and autonomous systems, ensuring both adversarial robustness uncertainty quantification is paramount. Conformal Prediction (CP) [4, 5, 6] offers a robust framework for uncertainty quantification by generating prediction sets that contain the true label at a predefined confidence level [7].
@@
The interplay between adversarial robustness and conformal prediction [8, 9] has received growing attention; however, its implications on the efficiency and robustness of CP sets remain underexplored. Previous studies have focused on adapting scoring functions when either test or calibration data are attacked, but these approaches often suffer from validity issues, undermining the robustness of CP sets under adversarial conditions.
In this work, we propose a novel integration of game-theoretic principles with conformal prediction to construct robust CP sets. Although our empirical evaluation is demonstrated using medical imaging datasets, the underlying framework is applicable to a wide range of domains. Our methodology specifically addresses the following research questions:
  • How can we construct valid and efficient CP sets under known adversarial attacks?
  • How can we maintain coverage guarantees in the presence of a set of potential adversarial perturbations?
  • What are the optimal defensive strategies in a zero-sum game setting where an adversary selects the most potent attack?
To tackle these questions, we develop specialized models by partitioning the training data and tailoring them to distinct attack types. We then construct conservative prediction sets based on the highest estimated quantile from calibration data across different adversarial scenarios. This conservative strategy ensures that the CP sets maintain their validity regardless of the attack. Moreover, by formulating the interplay between the defender and the adversary as a zero-sum game, we rigorously analyze the worst-case scenario—namely, the the most adverse adversarial perturbation that maximally impacts prediction set size an adversary can select—and derive optimal mixed defense strategies that minimize prediction set ambiguity while preserving coverage.
By integrating game-theoretic reasoning with CP, our approach enhances the robustness of deep learning systems in adversarial environments. In addition to maintaining high coverage rates, it minimizes prediction set sizes, thereby improving the safety and decision-making quality in critical applications. While our experiments focus on medical imaging, the proposed framework is not limited to this domain and can be extended to other applications where robustness to adversarial perturbations is crucial, such as unmanned driving, financial security analysis and network security protection.
@@ 2.1. Adversarial Attack
However, poisoning attacks are more difficult to detect in deep learning models [25, 26]. Consequently, a wider range of methods has been developed to enhance model robustness against such attacks. For instance, improvements have been made by modifying the model architecture [27], designing robust loss functions [28, 29], and optimizing loss functions [30].
@@ 2.2. Conformal Prediction under Adversarial Attack
Uncertainty estimation is crucial for the robustness of deep learning models. Conformal Prediction (CP), introduced by Vovk et al. [7], provides distribution-free coverage guarantees but faces challenges when confronted with data poisoning and adversarial attacks. Gendler et al. [31] proposed RSCP, enhancing CP’s robustness through randomized smoothing by replacing the scoring function to defend against $\ell_{2}$-norm adversarial perturbations, though its formal guarantees remain limited. Yan et al. [32] introduced RSCP+, which adds a hyperparameter $\beta$ and Hoeffding bounds to correct Monte Carlo errors, while incorporating RCT and PTT techniques to improve efficiency. However, these modifications often result in overly conservative prediction sets and increased reliance on the calibration set.
@@
In contrast, Ghosh et al. [33] proposed PRCP, which approaches adversarial perturbations from a probabilistic standpoint by relaxing robustness guarantees under a predefined distribution. Cauchois et al. [34] addressed distributional shifts using an f-divergence constraint, though the resulting optimization is overly conservative [35]. Angelopoulos et al. [8] extended CP to control expected loss but did not provide algorithms for computing worst-case adversarial loss. Einbinder et al. [9] demonstrated that standard CP is robust to random label noise, though adversarial label perturbations were not explored. Furthermore, traditional defenses based on anomaly detection, loss function adjustments, and data clustering [11, 12, 13, 15, 18] struggle to safeguard CP’s unique properties, especially against sophisticated poisoning attacks targeting uncertainty mechanisms. Additionally, Ennadir et al. [36] investigated the application of CP in adversarial attacks on graph neural networks.
@@
Building on these limitations, Liu et al. [37] proposed an uncertainty reduction-based adversarial training method (AT-UR), which combines Beta-weighted loss and entropy minimization to enable models to maintain robustness while generating larger and more meaningful prediction sets. Jeary et al. [38] introduced Verifiably Robust Conformal Prediction (VRCP), which supports perturbations bounded by various norms (including $\ell_{1}$, $\ell_{2}$, and $\ell_{\infty}$) and regression tasks, using recent neural network verification techniques to restore coverage guarantees under adversarial attacks.
On a different front, two distinct types of attacks have been explored in the context of Conformal Prediction (CP): evasion attacks and contamination (or poisoning) attacks. In evasion attacks, the adversary perturbs test inputs (often under norm constraints) to induce uncertainty or misclassification during inference. For instance, Zargarbashi et al. [39] proposed using the cumulative distribution function (CDF) of smoothed scores to derive tighter upper bounds on worst-case score variations. Their method addresses both continuous and discrete/sparse data by providing guarantees for evasion attacks (which perturb test inputs) as well as for contamination attacks (which perturb calibration data). In contrast, Li et al. [40] developed a new class of black-box data poisoning attacks that specifically target CP by manipulating the prediction uncertainty rather than directly causing misclassifications. They also proposed an optimization framework to defend against these poisoning attacks. Furthermore, Scholten et al. [41] introduced Reliable Prediction Sets (RPS), which aggregate predictions from classifiers trained on different partitions of the training data and calibrated on disjoint subsets, thereby enhancing CP’s resilience to contamination attacks.
@@
Recent studies have also highlighted CP’s resilience to label noise. Penso and Goldberger [42] introduced Noise-Robust Conformal Prediction (NR-CP), which estimates noise-free scores from noisy data, constructing smaller and more efficient prediction sets while maintaining coverage guarantees. NR-CP significantly outperforms other methods in noisy environments. Another study [9] explored the inherent robustness of CP, showing that even with noisy labels, CP tends to conservatively maintain coverage, although some adversarial noise conditions may compromise its robustness. Both studies emphasize CP’s resilience to label noise, a perturbation type that, while not intentionally malicious, introduces variability similar to adversarial attacks.
@@
In summary, our research complements and extends existing defense strategies by integrating game theory to model adversarial interactions and proposing novel mechanisms to fortify CP systems. This approach not only builds on the strengths of previous studies [33, 38, 40, 39] but also addresses their limitations, providing enhanced robustness for CP in high-stakes, adversarial environments.
@@ 3. Preliminary and Problem Setup
In this section, we establish the foundational notation and problem setup for evaluating and enhancing the robustness of CP systems against adversarial attacks. By defining the adversarial attack strategies, defensive models, conformal prediction framework, score functions, and metrics, and by formulating key research questions, we set the stage for developing and analyzing robust prediction mechanisms that maintain valid uncertainty estimates in the presence of deliberate perturbations. The notation used throughout the paper is given in Table 1.
@@ 3.1.1. Adversarial Attacks
Carlini &\&& Wagner (CW) Attack
The CW attack [47] is an optimization-based method for generating minimally perturbed adversarial examples. For the $L_{2}$ norm, the attack solves
$\min_{\delta}\;\|\delta\|_{2}^{2}+c\cdot f(x_{i}+\delta),$
subject to
$x_{i}+\delta\in[0,1]^{n},$
where $c>0$ balances the perturbation size with inducing misclassification, and $f$ is a loss function that encourages the attack. A typical definition for a targeted attack is
$f(x_{i}^{\prime})=\max\Big{\{}\max_{j\neq t}\{Z(x_{i}^{\prime})_{j}\}-Z(x_{i}^{\prime})_{t},-\kappa\Big{\}},$
with $t$ as the target class, $Z(x_{i}^{\prime})$ representing the model logits, and $\kappa$ controlling the confidence level. In practice, a change-of-variable such as
$x_{i}^{\prime}=\frac{1}{2}\Big{(}\tanh(w)+1\Big{)}$
is often adopted to enforce the input box constraint, where $w$ is the unconstrained variable optimized during the attack.
The CW attack is renowned for its effectiveness in producing adversarial examples that remain minimally perturbed while successfully causing misclassification, even on models that have been adversarially trained. Its optimization-based framework often makes it more challenging to defend against compared to simpler gradient-based methods.
Summary and Taxonomy of Attacks
The aforementioned adversarial attacks can be broadly categorized into several groups:
  • Gradient-based attacks: Including FGSM and PGD, these methods rely on the gradient information of the loss function. FGSM provides a one-step approximation, while PGD adopts an iterative procedure for improved attack strength.
  • Gradient-free attacks: Exemplified by SPSA, these approaches estimate gradients through random perturbations, making them suitable for scenarios where the model’s internal parameters are inaccessible.
  • Optimization-based attacks: Such as the CW attack, which formulates the attack as an optimization problem to achieve minimally perturbed yet effective adversarial examples.
In practice, an adversary may even consider using clean (i.e., unperturbed) data or selectively choose one of these attack methods depending on the defensive measures in place, thereby providing a comprehensive evaluation of model robustness against various adversarial perturbations. This selection process, where the adversary chooses an attack method that maximizes its impact on the prediction outcome (i.e., enlarging the prediction set) against a specific defense model, forms a core component of our game-theoretic analysis. Here, the attacker’s ability to choose among diverse attack strategies is explicitly modeled, reflecting the adaptive nature of adversarial threats.
@@ 3.3. Problem Formulation
To formalize these questions, consider the following setup. Let $f_{0}$ be the model trained on a clean training dataset $\{(x_{i},y_{i})\}_{i\in\mathcal{I}_{\textrm{train}}}$. An adversary leverages the available training labels to build an adversarial model and then applies an attack function $g_{j}$ on the test inputs $\{x_{i}\}_{i\in\mathcal{I}_{\textrm{test}}}$ (with the desired outcome known to the adversary) to generate perturbed inputs: $x^{\prime}_{i}=g_{j}(x_{i},y_{i},f_{0}).$ Note that while the notation includes $y_{i}$ for clarity in describing the mechanism, it is not necessary for the adversary to have access to the true labels of the test dataset. The objects from the test dataset, in conjunction with the adversary’s target outcome, are sufficient for carrying out the attack. Defensive models $\{f_{j}\}$ are then trained on the modified training set: $\{(x^{\prime}_{i},y_{i})\}_{i\in\mathcal{I}_{\textrm{train}}}$, so that each $f_{j}$ is tailored to resist the corresponding attack $g_{j}$.
@@ 4.2. Addressing RQ2: Unknown Adversarial Attack
1:Training dataset $\{(x_{i},y_{i})\}_{i\in\mathcal{I}_{\textrm{train}}}$, Calibration dataset $\{(x_{i},y_{i})\}_{i\in\mathcal{I}_{\text{cal}}}$, Post-unknown attack test dataset $\{x_{i}\}_{i\in\mathcal{I}_{\textrm{test}}}$, Set of potential attack functions $g_{j},j=1,\dots,m$, Coverage probability $1-\alpha$. 2:Prediction sets $\left(\Gamma(x_{i})\right)_{i\in\mathcal{I}_{\textrm{test}}}$ for a single defensive model $f_{k}$ 3:Based on the training method for $f_{k}$ outlined in Algorithm 2, we obtain the pre-trained defensive models $f_{k}$, $k=1,\dots,n$. $\triangleright$ Determine the largest quantile score across various attacks $g_{j}$ on the calibration set: 4:for each attack function $g_{j},j\in 1,\dots,m$ do       Apply attack $g_{j}$ to the calibration set to obtain:       $x^{\prime}_{i}=g_{j}(x_{i},y_{i},f_{k}),i\in\mathcal{I}_{\text{cal}}$.
      Use $f_{k}(x^{\prime}_{i})$ to compute non-conformity scores $\{s^{j}_{k}(x^{\prime}_{i},y_{i})\}_{i\in\mathcal{I}_{\text{cal}}}$.       Compute the quantile threshold as:       $q^{k_{j}}_{1-\alpha}$ as the $\lceil(1+|\mathcal{I}_{\text{cal}}|)(1-\alpha)\rceil$-th smallest score in ${s^{j}_{k}(x^{\prime}_{i},y_{i})}_{i\in\mathcal{I}_{\text{cal}}}$.
5:Determine the maximum quantile threshold $q^{k}_{1-\alpha}=\max_{j=1,\dots,m}q^{k_{j}}_{1-\alpha}$. $\triangleright$ Construct conservative prediction sets using the largest quantile score to ensure coverage: 6:for each test instance $x_{i}$, $i\in\mathcal{I}_{\textrm{test}}$ do       Construct the conservative prediction set for the defensive model $f_{k}(x^{\prime}_{i})$: $\Gamma(x_{i})=\{y\in\mathcal{Y}\mid s(x_{i},y)\leq q^{k}_{1-\alpha}\}$
Algorithm 3 Constructing Conformal Prediction Sets Across All Potential Attacks (RQ2)